# Coherent Perfect Absorption of Arbitrary Wavefronts at an Exceptional Point

### 05 Apr 2024 v1

@@ Abstract
(i) The first challenge is a CPA’s narrow bandwidth, with slight frequency shifts disrupting the resonant absorption, typically resulting in a Lorentzian absorption profile. Research on how to make CPAs more broadband encompasses a wide array of approaches, such as metasurfaces [27, 28, 29], plasmonic effects [30, 31, 32], metallic nanoparticles metasurfaces [33], thin films [34], and broadband PCN resonantors [35]. Recently also insights from Exceptional Point (EP) physics have been used for advancing this goal [11, 12, 13]: At an EP, two (or more) discrete eigenvalues and eigenstates of a non-Hermitian system merge. Such an EP-CPA can be observed when two (or more) cavities with spectrally-overlapping resonances are critically coupled as depicted in Fig. 1(b). Specifically, it is possible to select system parameters such that both the critical-coupling condition, neccessary for perfect absorption, and the EP-condition coincide at the same real input frequency [11]. The result is a spectral degeneracy that leads to a significantly broadened (quartic) absorption spectrum [12, 13], in analogy to the physics of white-light cavities [36, 37, 38, 39, 40, 41] and of multi-mirror Fabry-Pérot interferometers [42].
@@
Spatial and Spectral Degeneracy. — To derive the reflection behavior of light incident from the left on this MAD-EP-CPA we employ a scalar optics model for polarized light, considering a large, yet finite, number of spatial modes. Let $\textbf{R}_{1}$, $\textbf{R}_{2}$, and $\textbf{R}_{\mathrm{c}}$ represent the reflection matrices, and $\textbf{T}_{1}$, $\textbf{T}_{2}$, and $\textbf{T}_{\mathrm{c}}$ the transmission matrices for the left, right and central mirror, respectively. Additionally, $\textbf{T}_{\mathrm{a}}$ is the transmission through the absorber and $\textbf{T}_{\text{4f}}$ represents the transmission matrix for propagation through a $4f$-system, comprising two lenses in a telescopic configuration (without mirrors or absorbers). $\textbf{T}_{\text{4f}}$ acts as a double Fourier transform producing a spatially flipped image and a phase shift $e^{i\phi}$ that is the same for all modes. Using that the same reflection and transmission coefficients apply to all modes at the mirrors and the absorber, all matrices except $\textbf{T}_{\text{4f}}$ become diagonal matrices, e.g., $\textbf{R}_{1}=r_{1}\mathbb{1}$, $\textbf{T}_{1}=t_{1}\mathbb{1}$, etc. Using a transfer matrix approach, we can construct from these individual transmission matrices the sought-after reflection matrix $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(A)}}}$ for the entire MAD-EP-CPA [44, S1] (Design A in Fig. 1(d)):
@@
A detailed derivation is available in the supplementary material [44, S1-S3]. The result (2) shows that the relationship between $T_{\mathrm{a}}$ and $R_{1}$ for a MAD-EP-CPA is identical to that of a single-cavity MAD-CPA [43]: The higher the reflectivity $R_{1}$ of the input-coupling mirror, the weaker the internal absorber needs to be. Additionally, the relation between $R_{\text{c}}$ and $R_{1}$ indictates that $R_{\text{c}}$ quickly approaches values near $1$ with increasing values of $R_{1}$, indicating weak coupling between the left and right sub-cavity.
Numerical Simulations. — To corroborate that these analytical results are, indeed, applicable to the multi-mode MAD-EP-CPA, and to assess the impact of deviations in certain system parameters (such as minor misalignments of the mirrors), we conducted a corresponding numerical Scalar Fourier Optics computer simulation. As input we used a random speckle field, see Fig. 2(a), generated by 160 modes of equal amplitude and random phase in $k$-space. Without compensating for the refractive index of the absorber (which we assumed to be $n_{r}=1.5$, at a thickness of $d=0.6\mathrm{mm}$), our numerical simulation already indicates a considerably broadened absorption lineshape, but with the minimum reaching only 0.7%, see the red line in Fig. 2(b). To address this, we adjusted the focal length of the rightmost lens in Fig. 1(d) to $f^{\prime}=f-d(n_{\text{r}}-1/n_{\text{r}})/2$. This adjustment compensates for the increased optical path-length caused by the absorber and preserves identical optical path-lengths in both the left and right sub-cavity. As shown in the numerically calculated reflection spectrum of Fig. 2(b) (blue line), this correction restores perfect absorption at the point of degeneracy and verifies the successful operation of our MAD-EP-CPA cavity design (see the broad yellow line in Fig. 2(b) for comparison with an ideal quartic absorption lineshape). For a more detailed explanation of this refraction compensation technique, readers are referred to the supplementary text [44, S8]. Another potential strategy to minimize refractive aberrations is to use an absorber with a refractive index whose real part is as close to $1$ as possible [45].
@@
The numerical simulations also allow us to study the impact of small deviations from optimal system parameters, and help to pinpoint the critical system parameters influencing the system performance. For instance, the simulations demonstrate that the system exhibits robustness to deviations of the optimal reflectivity $R_{1}$ of the input coupling mirror (to within about ±$\pm 5\%$), or of the optimal transmissivity $T_{\mathrm{a}}$ of the absorber (to within about ±$\pm 2\%$). In contrast, deviations in the reflectivity $R_{\mathrm{c}}$ of the central coupling mirror from its optimal value significantly affect the system performance. Even a mere $1\%$ deviation from the optimal value results in a minimum reflectivity of approximately $5\%$, significantly diminishing the system’s absorption efficiency (further details and figures can be found in the supplementary text [44, S4]). As expected, the precise parallel alignment of the mirrors is also crucial, particularly for the central coupling mirror, as shown in Fig. 3. Additional details on critical parameters, such as the consequences of slight discrepancies in the alignment of the two cavity lengths or of residual lens reflections, are available in the supplementary text [44, S4].
We have also studied MAD-EP-CPA systems encompassing more than two coupled cavities, as detailed in the supplementary text [44, S5]. Interestingly, constructing an EP-CPA with an increasing number of cavities results in a reflection spectrum around the resonance frequencies akin to a Butterworth bandpass filter, a concept from electrical filter theory [46, 47]. This observation is consistent with the established understanding that, as outlined in [42], the Butterworth function also characterizes the transmission behavior of a multi-mirror Fabry-Pérot interferometer in the absence of a dissipative element.
@@
However, Design A poses challenges in the implementation, as it involves weakly coupling two plane-parallel cavities with altogether four embedded lenses. We thus propose in Fig. 4(a) an alternative “Design B”, involving a coupling mirror with reflectivity $R_{\mathrm{c}}$ at the center of a single $4f$-cavity, requiring only two lenses. This simplification requires a division of the left mirror at the optical axis into an input-coupling segment with reflectivity $R_{1\mathrm{a}}<1$ and a fully reflective mirror $R_{1\mathrm{b}}=1$. The weak absorber is then placed adjacent to the totally reflective mirror $R_{1\mathrm{b}}$. Additionally, to mitigate the additional refraction introduced by the absorber, we adjust the focal length of the left lens to $f^{\prime}=f-d(n_{\text{r}}-1/n_{\text{r}})/2$ (see supplementary [44, S8]). Furthermore, to ensure an identical optical path-length above and below the optical axis, we have incorporated a transparent slab, designated $T_{1}$, having the same thickness and refractive index as the absorber, adjacent to mirror $R_{1\mathrm{b}}$.
Design B can be seen as a “folded-up” variant of Design A by identifying analogous round-trip paths. For instance, in Design A, light undergoes self-imaging round-trips within the right sub-cavity, with a corresponding path in Design B from the bottom-left to the top-right. A similar parallel can be observed for round-trips in Design A’s left sub-cavity, corresponding with the path from top-left to bottom-right in Design B. In both of these scenarios, a single reflection at Design A’s central mirror is substituted by two transmissions through Design B’s central mirror. Also, for the third round-trip-path in Design A, where light propagates between the farthest mirrors $R_{1}$ and $R_{2}$, an analog round-trip-path can be identified in Design B (see supplementary [44, S7] for details). However, it is crucial to recognize that such comparisons between Design A and Design B offer, at best, a qualitative analogy. A comprehensive comparison requires a sound mathematical analysis, considering the infinite sum of all possible round trips and the effects of coupling between the two sub-cavities. Nonetheless, this qualitative comparison hints that the condition for the transmissivity $T_{\text{c}}$ of the central mirror in Design B may resemble the condition for the reflectivity $R_{\text{c}}$ of the central mirror in Design A. Using an analogous derivation method as before, the following matrix equation for $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(B)}}}$ can be established:
@@
with the scalar value $a=r_{\text{c}}^{2}(t_{\text{a}}^{2}+1)(r_{\text{1a}}^{2}+1)$. Observe how the structure of Eq. (3) resembles the structure of Eq. (1). Because $\textbf{T}_{\text{2f}}^{4}=\textbf{T}_{\text{4f}}^{2}$, the matrix $\textbf{T}_{\text{2f}}^{4}$ in Eq. (3) results in a quadruple Fourier transformation with a uniform phase shift $e^{2i\phi}$ for all modes, and can thus be represented as a simple diagonal matrix $\textbf{T}_{\text{4f}}^{2}=e^{2i\phi}\mathbb{1}$. Analogously, $\textbf{T}_{\text{2f}}^{8}=e^{4i\phi}\mathbb{1}$. Consequently, $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(B)}}}$ becomes a diagonal matrix and can be reduced to a simple scalar equation, which proves the spatial degeneracy of Design B. The spectral degeneracy condition can be met when the following EP-conditions are fulfilled (see supplementary [44, S6]):

### 08 Apr 2024 v2

@@ Abstract
(i) The first challenge is a CPA’s narrow bandwidth, with slight frequency shifts disrupting the resonant absorption, typically resulting in a Lorentzian absorption profile. Research on how to make CPAs more broadband encompasses a wide array of approaches, such as metasurfaces [27, 28, 29], plasmonic effects [30, 31, 32], metallic nanoparticles metasurfaces [33], thin films [34], and broadband PCN resonators [35]. Recently also insights from Exceptional Point (EP) physics have been used for advancing this goal [11, 12, 13, 36]: At an EP, two (or more) discrete eigenvalues and eigenstates of a non-Hermitian system merge. Such an EP-CPA can be observed when two (or more) cavities with spectrally-overlapping resonances are critically coupled as depicted in Fig. 1(b). Specifically, it is possible to select system parameters such that both the critical-coupling condition, neccessary for perfect absorption, and the EP-condition coincide at the same real input frequency [11]. The result is a spectral degeneracy that leads to a significantly broadened (quartic) absorption spectrum [12, 13], in analogy to the physics of white-light cavities [37, 38, 39, 40, 41, 42] and of multi-mirror Fabry-Pérot interferometers [43].
@@
Spatial and Spectral Degeneracy. — To derive the reflection behavior of light incident from the left on this MAD-EP-CPA we employ a scalar optics model for polarized light, considering a large, yet finite, number of spatial modes. Let $\textbf{R}_{1}$, $\textbf{R}_{2}$, and $\textbf{R}_{\mathrm{c}}$ represent the reflection matrices, and $\textbf{T}_{1}$, $\textbf{T}_{2}$, and $\textbf{T}_{\mathrm{c}}$ the transmission matrices for the left, right and central mirror, respectively. Additionally, $\textbf{T}_{\mathrm{a}}$ is the transmission through the absorber and $\textbf{T}_{\text{4f}}$ represents the transmission matrix for propagation through a $4f$-system, comprising two lenses in a telescopic configuration (without mirrors or absorbers). $\textbf{T}_{\text{4f}}$ acts as a double Fourier transform producing a spatially flipped image and a phase shift $e^{i\phi}$ that is the same for all modes. Using that the same reflection and transmission coefficients apply to all modes at the mirrors and the absorber, all matrices except $\textbf{T}_{\text{4f}}$ become diagonal matrices, e.g., $\textbf{R}_{1}=r_{1}\mathbb{1}$, $\textbf{T}_{1}=t_{1}\mathbb{1}$, etc. Using a transfer matrix approach, we can construct from these individual transmission matrices the sought-after reflection matrix $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(A)}}}$ for the entire MAD-EP-CPA [45, S1] (Design A in Fig. 1(d)):
@@
A detailed derivation is available in the supplementary material [45, S1-S3]. The result (2) shows that the relationship between $T_{\mathrm{a}}$ and $R_{1}$ for a MAD-EP-CPA is identical to that of a single-cavity MAD-CPA [44]: The higher the reflectivity $R_{1}$ of the input-coupling mirror, the weaker the internal absorber needs to be. Additionally, the relation between $R_{\text{c}}$ and $R_{1}$ indictates that $R_{\text{c}}$ quickly approaches values near $1$ with increasing values of $R_{1}$, indicating weak coupling between the left and right sub-cavity.
Numerical Simulations. — To corroborate that these analytical results are, indeed, applicable to the multi-mode MAD-EP-CPA, and to assess the impact of deviations in certain system parameters (such as minor misalignments of the mirrors), we conducted a corresponding numerical Scalar Fourier Optics computer simulation. As input we used a random speckle field, see Fig. 2(a), generated by 160 modes of equal amplitude and random phase in $k$-space. Without compensating for the refractive index of the absorber (which we assumed to be $n_{r}=1.5$, at a thickness of $d=0.6\mathrm{mm}$), our numerical simulation already indicates a considerably broadened absorption lineshape, but with the minimum reaching only 0.7%, see the red line in Fig. 2(b). To address this, we adjusted the focal length of the rightmost lens in Fig. 1(d) to $f^{\prime}=f-d(n_{\text{r}}-1/n_{\text{r}})/2$. This adjustment compensates for the increased optical path-length caused by the absorber and preserves identical optical path-lengths in both the left and right sub-cavity. As shown in the numerically calculated reflection spectrum of Fig. 2(b) (blue line), this correction restores perfect absorption at the point of degeneracy and verifies the successful operation of our MAD-EP-CPA cavity design (see the broad yellow line in Fig. 2(b) for comparison with an ideal quartic absorption lineshape). For a more detailed explanation of this refraction compensation technique, readers are referred to the supplementary text [45, S8]. Another potential strategy to minimize refractive aberrations is to use an absorber with a refractive index whose real part is as close to $1$ as possible [46].
@@
The numerical simulations also allow us to study the impact of small deviations from optimal system parameters, and help to pinpoint the critical system parameters influencing the system performance. For instance, the simulations demonstrate that the system exhibits robustness to deviations of the optimal reflectivity $R_{1}$ of the input coupling mirror (to within about ±$\pm 5\%$), or of the optimal transmissivity $T_{\mathrm{a}}$ of the absorber (to within about ±$\pm 2\%$). In contrast, deviations in the reflectivity $R_{\mathrm{c}}$ of the central coupling mirror from its optimal value significantly affect the system performance. Even a mere $1\%$ deviation from the optimal value results in a minimum reflectivity of approximately $5\%$, significantly diminishing the system’s absorption efficiency (further details and figures can be found in the supplementary text [45, S4]). As expected, the precise parallel alignment of the mirrors is also crucial, particularly for the central coupling mirror, as shown in Fig. 3. Additional details on critical parameters, such as the consequences of slight discrepancies in the alignment of the two cavity lengths or of residual lens reflections, are available in the supplementary text [45, S4].
We have also studied MAD-EP-CPA systems encompassing more than two coupled cavities, as detailed in the supplementary text [45, S5]. Interestingly, constructing an EP-CPA with an increasing number of cavities results in a reflection spectrum around the resonance frequencies akin to a Butterworth bandpass filter, a concept from electrical filter theory [47, 48]. This observation is consistent with the established understanding that, as outlined in [43], the Butterworth function also characterizes the transmission behavior of a multi-mirror Fabry-Pérot interferometer in the absence of a dissipative element.
@@
However, Design A poses challenges in the implementation, as it involves weakly coupling two plane-parallel cavities with altogether four embedded lenses. We thus propose in Fig. 4(a) an alternative “Design B”, involving a coupling mirror with reflectivity $R_{\mathrm{c}}$ at the center of a single $4f$-cavity, requiring only two lenses. This simplification requires a division of the left mirror at the optical axis into an input-coupling segment with reflectivity $R_{1\mathrm{a}}<1$ and a fully reflective mirror $R_{1\mathrm{b}}=1$. The weak absorber is then placed adjacent to the totally reflective mirror $R_{1\mathrm{b}}$. Additionally, to mitigate the additional refraction introduced by the absorber, we adjust the focal length of the left lens to $f^{\prime}=f-d(n_{\text{r}}-1/n_{\text{r}})/2$ (see supplementary [45, S8]). Furthermore, to ensure an identical optical path-length above and below the optical axis, we have incorporated a transparent slab, designated $T_{1}$, having the same thickness and refractive index as the absorber, adjacent to mirror $R_{1\mathrm{b}}$.
Design B can be seen as a “folded-up” variant of Design A by identifying analogous round-trip paths. For instance, in Design A, light undergoes self-imaging round-trips within the right sub-cavity, with a corresponding path in Design B from the bottom-left to the top-right. A similar parallel can be observed for round-trips in Design A’s left sub-cavity, corresponding with the path from top-left to bottom-right in Design B. In both of these scenarios, a single reflection at Design A’s central mirror is substituted by two transmissions through Design B’s central mirror. Also, for the third round-trip-path in Design A, where light propagates between the farthest mirrors $R_{1}$ and $R_{2}$, an analog round-trip-path can be identified in Design B (see supplementary [45, S7] for details). However, it is crucial to recognize that such comparisons between Design A and Design B offer, at best, a qualitative analogy. A comprehensive comparison requires a sound mathematical analysis, considering the infinite sum of all possible round trips and the effects of coupling between the two sub-cavities. Nonetheless, this qualitative comparison hints that the condition for the transmissivity $T_{\text{c}}$ of the central mirror in Design B may resemble the condition for the reflectivity $R_{\text{c}}$ of the central mirror in Design A. Using an analogous derivation method as before, the following matrix equation for $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(B)}}}$ can be established:
@@
with the scalar value $a=r_{\text{c}}^{2}(t_{\text{a}}^{2}+1)(r_{\text{1a}}^{2}+1)$. Observe how the structure of Eq. (3) resembles the structure of Eq. (1). Because $\textbf{T}_{\text{2f}}^{4}=\textbf{T}_{\text{4f}}^{2}$, the matrix $\textbf{T}_{\text{2f}}^{4}$ in Eq. (3) results in a quadruple Fourier transformation with a uniform phase shift $e^{2i\phi}$ for all modes, and can thus be represented as a simple diagonal matrix $\textbf{T}_{\text{4f}}^{2}=e^{2i\phi}\mathbb{1}$. Analogously, $\textbf{T}_{\text{2f}}^{8}=e^{4i\phi}\mathbb{1}$. Consequently, $\textbf{R}_{\text{\tiny{CPA}}}^{\text{\tiny{(B)}}}$ becomes a diagonal matrix and can be reduced to a simple scalar equation, which proves the spatial degeneracy of Design B. The spectral degeneracy condition can be met when the following EP-conditions are fulfilled (see supplementary [45, S6]):